in the folloeing reaction xA=yB log[-d[A]/dt]=log[d[B]/dt]+0.3 thus x:y ratio is:a)1:2b)2:1c)3:1d)3:10

Grade:11

1 Answers

sam
6 years ago
it will be 2:1 as we know for a simple chemical reaction
average rate of dissapearance= average rate of apperance
A → B
-dA/dT= dB/dT  
So for this reaction, (-dA/dT)/ x = (dB/dT)/y                                          ….(i)
We can write 0.3 as log 2 (Approx) 
then applying property of log... log a + log b= log ab
we get log(-dA/dT)= log 2(dB/dT)
taking antilog both sides and dividing by 2 and compare with (i)
x=2, y=1 hence x:y =2:1